An electrically driven road vehicle of mass 450kg has four wheels of effective diameter 400mm radius of gyration 125mm and mass 9kg. The armature of the electric motor has a mass of 65kg a radius of gyration of 100mm and rotates at 4 times the speed of the road wheels. The rolling resistance to motion of the vehicle is to be assumed constant at 225N. Determine the acceleration of the vehicle if it is allowed to run freely down a slope whose inclination to the horizontal is sin-¹ 0.2


1
Expert's answer
2021-09-21T02:06:49-0400

mv=450kg

dw=0.4m

kw=0.125m

mw=9kg

ma=65kg

ka=0.1m

Rotates 4 times speed of road wheels

Fr=225N

a=?

From energy balance;

rate of P.E loss of car = rate of rotational K.E gain of armature + rate of rotational K.E gain of wheels + rate of K.E gain of car - rate of friction dissipation

Rate of P.E loss of car = mv*g*vSin"\theta"

Rate of rotational K.E gain of armature = Ia*wa*wa'

Rate of rotational K.E gain of wheels = Iw*w*w'

Rate of K.E gain of car = mv*v*v''

Rate of friction dissipation = F*v

0.2*mv*g*v=64*Ia*(v)"\frac{(a)}{d^{2}}" +4*Iw*(v)"\frac{(a)}{d^{2}}" +mv*v*a-F*v

0.2*mv*g=64*Ia*"\frac{(a)}{d^{2}}" +4*Ia*"\frac{(a)}{d^{2}}" +mv*a-F

0.2*450*9.81=a(64*65*"\frac{0.1^{2}}{0.4^{2}}" +450)-225

a="\frac{1107.9}{710}"

="1.5604m\/s^{2}"